### v1

# High-order regularization dealing with ill-conditioned robot localization problems

Abstract
In this work, we propose a high-order regularization method to solve the ill-conditioned problems in robot localization. Numerical solutions to robot localization problems are often unstable when the problems are ill-conditioned. A typical way to solve ill-conditioned problems is regularization, and a classical regularization method is the Tikhonov regularization. It is shown that the Tikhonov regularization can be seen as a low-order case of our method. We find that the proposed method is superior to the Tikhonov regularization in approximating some ill-conditioned inverse problems, such as robot localization problems. The proposed method overcomes the over-smoothing problem in the Tikhonov regularization as it can use more than one term in the approximation of the matrix inverse, and an explanation for the over-smoothing of the Tikhonov regularization is given. Moreover, one a priori criterion which improves the numerical stability of the ill-conditioned problem is proposed to obtain an optimal regularization matrix. As most of the regularization solutions are biased, we also provide two bias-correction techniques for the proposed high-order regularization. The simulation and experiment results using a sensor network in a 3D environment are discussed, demonstrating the performance of the proposed method.
Index Terms:
Robot localization, ill-conditioned problem, high-order regularization, approximate solution, a priori criterion for numerical stability.
I Introduction
The growing usage of mobile robots, e.g. indoor parking lots, cargo automation management in automatic factories, autonomous vehicles, and fire rescue [1], gives rise to the high demand of localization techniques, especially under limited or constrained sensing capabilities. A challenging problem in robot localization is to deal with the ill-conditioned situations[2], in which small changes in the measurements lead to huge variations in the location solutions [3]. Regularization introduces penalty terms when solving for location solutions and is one of the most direct approaches preventing the appearance of ill conditions. The most typical regularization method is Tikhonov regularization (TR) [4] (sometimes referred to, under generalization, as the $L_{2}$-norm regularization [5]). An improved TR is proposed by Fuhry et al. [6], which shows the relation between the truncated singular value decomposition (TSVD) method and the TR method.
The improved TR provides a special choice of the regularization matrix, which is further developed by Yang et al. [7], Noschese et al.[8], Mohammady et al.[9] and Cui et al.[10]. A fractional filter method [11] is proposed by Klann and developed by Mekoth who presents the fractional TR method [12]. A thorough overview of modern regularization methods is covered in [4]. For a regularization method, such as the TR method, one of the key procedures is the determination of regularization parameters. There are many methods to select the regularization parameter, such as L-curve and generalized cross-validation (GCV) methods [13]. However, these parameter-selecting methods are posterior or empirical methods, which means that to acquire the regularization parameters, iterative algorithms or data related to the specific problems and environments are required. Therefore, they are computationally expensive [14]. The bias of the regularization solution is also a main topic in regularization. Bias-correction methods using iteratively estimated variance components, proposed by Xu et al.[15], Shen et al. [16] and Ji et al.[17], are proven to be effective to calculate the bias and achieve unbias solutions for some problems[16].
Applying and developing regularization to robot localization problems and their related areas in robotics have been an intensively studied area over the past decade [23, 18, 19, 20, 22, 21]. Nerurkar et al. apply a Tikhonov regularization-based algorithm, the Levenberg-Marquardt (LM) algorithm, to estimate robots’ poses in cooperative robot localization problems[23]. Wang et al. [18] solve an ill-conditioned linearized localization problem using the TR method. Yuan et al. [19] present a factor graph-based framework and a regularized least square (RLS) for passive localization problems. Ning et al. [22] study a localization problem with truncated singular value decomposition regularization. Liouane et al. [20] improve the location estimation accuracy of the sensor in a wireless sensor network with a regularized least square method. Pan et al. [21] work on applying TR to the multi-sensor localization problem. Tang et al. [24] and Hu et al. [25] study the underwater localization problems with regularization to deal with the ill-conditioned situations. Details of the Tikhonov regularization-based methods are studied [26, 27, 28] to solve specific problems in related areas of robot localization. The Tikhonov regularization is introduced into a kernel learning method to provide the generalization ability of a kernel model and improve the localization behavior of the kernel model when the robot is in high dynamics conditions[26]. Yu et al.[27] study the three-dimensional passive localization problem that is vulnerable to the influence of measurement error and apply TR to solve the ill-conditioned localization problem. Fang et al. [28] propose a robust Newton algorithm based on TR by avoiding the ill-conditioned Hessian matrix caused by the bad initial value. Their results show improvements in the accuracy, stability, and convergence performance of the proposed algorithms. Other applications of Tikhonov regularization include path-planning for small-scale robots [29], motion control of a magnetic microrobot [30], and calibration of a robot positioning processes [31]. While Han et al. work on applying $L_{1}$ regularization for collision avoidance in safe robot navigation [32], Wu et al. focus on the analysis of the optimal TR matrix and its application for localization problems with global navigation satellite system (GNSS) [33]. However, calculating the proposed optimal regularization matrix requires prior information (the first and second central moments) of an unknown integer ambiguity vector. Some regularization methods, e.g., Bayesian regularization, which can be simplified as Tikhonov regularization in some cases [34, 35], are used to train deep neural networks (DNNs) using the LM algorithm and then the DNNs are used to localize indoor robots to achieve high positioning precision[36]. Regularization in safe and smooth robot localization when considering continuous movements of robots[37] is studied by Dümbgen. Dong et al. [38] study a localization problem of a flying robot with a single UWB ranging measurement and IMU measurement, in which they introduce velocity penalty terms to the objective function of their problem and solve the problem by the LM method with regularization. The use of regularization in vision tasks, such as feature detection and matching, 3-D point cloud registration, and visual homing, which can be applied in simultaneous localization and mapping (SLAM) for robot localization problems [39, 40, 41], is also studied as there are many ill-conditioned situations in SLAM. However, additional efforts need to be made to analyze the impact and explain the effect of regularization on the solutions.
This paper proposes a high-order regularization method to solve ill-conditioned problems and get stable numerical solutions to robot localization tasks. The main contributions of this work are summarized as follows.
  1. 1.
    A new high-order regularization method is constructed. The proposed high-order regularization method explains and overcomes the over-smoothing problem in the Tikhonov regularization. It also bridges the least square method and the Tikhonov regularization method mathematically by approximating the inverse of a matrix. According to the approximation expression of the matrix inverse, the lower bound and the upper bound of the approximation residual are provided, which is used to derive the error bounds of the high-order regularization solutions, as well as the error bounds of the Tikhonov regularization solutions as TR is proven to be a low-order case of high-order regularization.
  2. 2.
    We provide a new relaxation method for selecting the regularization matrix. The presented relaxation method is used to improve the numerical stability of the algorithm by modifying the condition number using the regularization matrix.
  3. 3.
    We construct one a priori criterion to obtain the optimal regularization matrix for the proposed high-order regularization method. The regularization matrix is explicitly and directly calculated according to the proposed criterion. The closed-form solution of the optimal regularization matrix under the criterion is also provided for some cases.
  4. 4.
    Two bias-correction methods are provided to drive a possible unbiased high-order regularization.
  5. 5.
    A bias-correction algorithm with a sliding window for the robot localization problems is proposed, and the ill-conditioned situation for the robot localization problem is discussed.
  6. 6.
    The implementation of the high-order regularization method and the bias-correction algorithm for robot localization in 3D environments are also presented.
The rest of this paper is organized as follows. In Section II, regularization for ill-conditioned problems in robot localization is introduced. In Section III, a high-order regularization method is proposed, which shows the relationship between the LS method and the TR method and the reason for the over-smoothness of the TR solution. The relaxation method and a prior criterion are proposed to obtain the optimal regularization matrix of the high-order regularization method. Two bias-correction techniques for the high-order regularization method are also presented in Section III. In Section IV, robot localization using a sensor network is discussed, and a bias-correction method with a sliding window is described. In Section V, some simulation results, as well as experimental results, are given to demonstrate the performance of the proposed method. The summarization and future work of this paper are provided in Section VI.
II Ill-conditioned problems in robot localization and their Tikhonov regularization
To solve ill-conditioned robot localization problems, we propose a high-order regularization method and a theoretical solution for selecting the regularization parameter based on a priori criterion. The proposed method is superior to the TR method and overcomes the over-smoothing problem in the TR method.
A range of localization problems, such as robot localization problems using a sensor network, Ultra-Wideband (UWB) aided robot localization, pose initialization in SLAM [42, 43, 44, 45, 46], can be described using linear models; even when the original models are nonlinear, their linearized approximation [47] can still lead to meaningful solutions. Consider a set of linear equations
$Ax=b$ (1)
or its related minimization problem
$\min_{x\in\mathbb{R}^{n}}\|Ax-b\|$ (2)
[…]
$\hat{x}_{ls}=\left(A^{T}A\right)^{-1}A^{T}b$ (3)
However, it is more challenging when the matrix $A$ is ill-conditioned ([48], Chapter 5.8); in this case, an approximate solution given by the TR method [49] can be handy. A standard form of the TR solution can be found in [49], which is given by
$\hat{x}_{tr}=\left(A^{T}A+\mu^{2}I\right)^{-1}A^{T}b$ (4)
where $\mu$ is a given regularization parameter. This solution solves the following problem
$\min_{x\in\mathbb{R}^{n}}\left\{\|Ax-b\|+\mu^{2}\|x\|\right\}$ (5)
which is an approximate problem of the original problem (2) with an extra penalty term called the regularization term.
It is known, though, that the TR solution can be overly smooth [11] without a proper justification of adding a regularization term to the original problem (2). We will explicitly show that the TR solution can be seen as an approximate solution in terms of matrix inverse approximation using the LS method by adding the regularization term. Computing the inverse of the ill-conditioned matrix is the main reason that the LS solution is unstable.
III New high-order regularization method
To improve the numerical stability of ill-conditioned problems, which, in other words, turns the ill-conditioned problems into well-conditioned ones, and get a numerically stable solution to the original problems (1), we propose a high-order regularization (HR) method. Assuming that $A$ is known exactly, the $k$th-order regularization solution is
$\hat{x}^{k}_{hr}=\left(A^{T}A+R\right)^{-1}\sum_{i=0}^{k}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}A^{T}b$ (6)
where the index $k$ is the order of the proposed method, the regularization matrix $R$ satisfies the spectral radius condition $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$, and the matrix $A^{T}A$ and the new matrix $\left(A^{T}A+R\right)$ are nonsingular. We will show that if $R\in\mathrm{S}_{+}^{n}$, the convex set of symmetric positive semidefinite (PSD) matrices, then the spectral radius condition has always been satisfied. And then the nonsingular matrix $\left(A^{T}A+R\right)\in\mathrm{S}_{++}^{n}$, which is the convex set of the symmetric positive definite (SPD) matrix. The equation (6) gives a $k$th-order regularization approximate solution to the equation (1). And then, to obtain an optimal regularization matrix, we introduce the a priori criterion with its loss function given as follows
$\min_{k\in\mathbb{N},R\in\mathbb{R}^{n\times n}}f_{0}(k,R)$ (7)
$=\min_{k\in\mathbb{N},R\in\mathbb{R}^{n\times n}}\begin{array}[]{l}\bigg{\{}\left\|\sum_{i=k+1}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}\right\|\\ \quad+\left\|\left(A^{T}A+R\right)\right\|\left\|\left(A^{T}A+R\right)^{-1}\right\|\bigg{\}}\\ \end{array}$
[…]
The proposed criterion (7) is a priori which can be determined by the known matrix $A$ before any estimation of the solution of the problem, taking into account the trade-off between the estimation bias of the high-order regularization method and the improvement in the numerical stability of the approximation solution of the problem (2) by using the new matrix $\left(A^{T}A+R\right)$ with a small condition number. And for any problem described by (2), the proposed criterion (7) can be used to find the new (SPD) matrix $\left(A^{T}A+R\right)$ with a small condition number to calculate the approximation solution of the problem.
In the following subsections, we present the details of the proposed high-order regularization methods.
III-A Matrix inverse and the high-order regularization method
To obtain the high-order regularization solution, we first recall the matrix series [48]
$\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}=\sum_{i=0}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i},$ (8)
[…]
Thus, the inverse of the matrix $A^{T}A$ is
$\left(A^{T}A\right)^{-1}=\left(\left(I-R\left(A^{T}A+R\right)^{-1}\right)\left(A^{T}A+R\right)\right)^{-1}$ (9)
$=\left(A^{T}A+R\right)^{-1}\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}$
$=\left(A^{T}A+R\right)^{-1}\sum_{i=0}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}$
since the matrix $A^{T}A$ can be rewritten as
$A^{T}A=\left(I-R\left(A^{T}A+R\right)^{-1}\right)\left(A^{T}A+R\right)\text{. }$ (10)
Considering the first $k+1$ terms of the matrix power series in equation (9), one has the $k$th-order regularization method (6).
Lemma 1.
The TR solution given in (4) is a special case of the proposed HR solution (6) for $k=0$.
Proof.
When $k=0$, the HR solution reduces to
$\hat{x}^{0}_{hr}=\left(A^{T}A+R\right)^{-1}A^{T}b$ (11)
which is equivalent to the TR solution with $R=\mu^{2}I$, and the spectral radius condition $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$ required in the matrix series can always be satisfied. ∎
In general, one can select $k>0$, and a regularization matrix that meets the spectral radius condition to alleviate the over-smoothing of the TR method. For $k=1$, the HR solution can be simplified to
$\hat{x}^{1}_{hr}=$ $\left(A^{T}A+R\right)^{-1}A^{T}b$ (12)
$+\left(A^{T}A+R\right)^{-1}R\left(A^{T}A+R\right)^{-1}A^{T}b$
$=$ $\left(I+\left(A^{T}A+R\right)^{-1}R\right)\left(A^{T}A+R\right)^{-1}A^{T}b$
which is a better solution than the TR solution of problem (2) in the sense of approximation of the inverse of the matrix $A^{T}A$. To alleviate the over-smoothness of the TR method, the equation (12) shows that adding some extra terms is equivalent to multiplying by an appropriate factor $\left(I+\left(A^{T}A+R\right)^{-1}R\right)$.
Lemma 2.
If $R$ is symmetric positive semidefinite and $A^{T}A$ is positive definite, then the spectral radius condition $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$ has always been satisfied.
Proof.
If $R$ is symmetric positive semidefinite and $A^{T}A$ is symmetric positive definite, then $\left(A^{T}A+R\right)$ is symmetric positive definite and $A^{T}A+R\succ R$, where $\succ$ defines the partial ordering of symmetric positive definite matrices [48]. And then according to [48] (Theorem 7.7.3, p. 494), it follows that the spectral radius condition is satisfied. ∎
Since one can always design a symmetric regularization matrix $R$ for robot localization problems, according to Lemma 2, the required spectral radius condition will always be satisfied in such a selection. We will introduce some strategies to determine an appropriate regularization matrix later in the paper.
Proposition 1.
Let $R\in\mathrm{S}_{+}^{n},A^{T}A\in\mathrm{S}_{++}^{n}$, then the regularization matrix $R$ and the inverse $(A^{T}A+R)^{-1}$ can be diagonalized simultaneously via similarity by some unitary matrix if and only if $R$ and $(A^{T}A+R)^{-1}$ commute; if and only if $R\left(A^{T}A+R\right)^{-1}$ is normal; and if and only if the matrix $R\left(A^{T}A+R\right)^{-1}$ is a positive semidefinite matrix.
Proof.
Take $H=\left(A^{T}A+R\right)^{-1}\in\mathrm{S}_{++}^{n}$, then $H$ can be diagonalized via similarity by some unitary matrix. According to [48] (Theorem 1.3.12, p. 62 and Theorem 4.1.6, p. 229), $R$ and $H$ can be diagonalized simultaneously by some unitary matrix if and only if $R$ and $H$ commute. From [48] (Theorem 4.5.15, p. 286) or [50] (Theorem 3), $R$ and $H$ commute if and only if $RH$ is normal, and if and only if $RH$ is a positive semidefinite matrix. ∎
III-B Explanation of regularization methods
From equation (9), one obtains an approximation expression of the inverse of $A^{T}A$. To approximate the inverse of the matrix $A^{T}A$, the TR method uses the regularization matrix $\mu^{2}I$ in its solution. As the TR solution only contains one term of the approximation expression (9), that is the main reason why the TR solution is often overly smooth.
In our proposed method, one can use more than one term of the approximation expression, which enhances the approximate performance of the HR solution. In general, using the first two terms ($k=1$) of the matrix series is a wise choice, which will not require too much extra computation.
One can improve the numerical stability of the ill-conditioned problem described by equation (1) by calculating its approximate solution using the new matrix $(A^{T}A+R)$ with a small condition number, which can be achieved by selecting a suitable regularization matrix $R$. It is known that matrix inversion is very computationally expensive in matrix operations. The computing result of $(A^{T}A+R)^{-1}$ can be reused in one calculation of the HR solution. This means that the computation expenses of the HR method and TR method are almost the same. And the approximation expression (9) also provides a fast and computable approximate solution for calculating the inverse of a matrix.
III-C Error bound
We discuss the error bound [12] of the HR method in this section. First, one can express a matrix series as follows
$F\left(R\left(A^{T}A+R\right)^{-1}\right):=\sum_{i=k+1}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i},$ (13)
where $k$ is the same index of the equation (6). Then the bias between the LS solution and the HR solution can be obtained by
$\hat{x}_{ls}-\hat{x}^{k}_{hr}=\left(A^{T}A+R\right)^{-1}F\left(R\left(A^{T}A+R\right)^{-1}\right)A^{T}b.$ (14)
The error bound of the estimate HR method is
$\left\|\hat{x}_{ls}-\hat{x}^{k}_{hr}\right\|\leq\left\|\left(A^{T}A+R\right)^{-1}F\left(R\left(A^{T}A+R\right)^{-1}\right)A^{T}b\right\|.$ (15)
When $k=1$, the bias has another similar expression according to a similar approximation expression of the matrix $A^{T}A$ in the LS solution
$\hat{x}_{ls}-\hat{x}^{1}_{hr}=\left(A^{T}A+R\right)^{-1}\sum_{i=2}^{\infty}\left(-R\left(A^{T}A\right)^{-1}\right)^{i}A^{T}b$ (16)
$=\left(A^{T}A+R\right)^{-1}$
×$\quad\times\left(\left(I+R\left(A^{T}A\right)^{-1}\right)^{-1}-I+R\left(A^{T}A\right)^{-1}\right)A^{T}b$
which requires the condition $\rho\left(R\left(A^{T}A\right)^{-1}\right)<1$ (see Appendix A).
The difference between the TR solution and the LS solution is
$\left\|\hat{x}_{tr}-\hat{x}_{ls}\right\|$ (17)
$\leq\left\|\left(A^{T}A+R\right)^{-1}\sum_{i=1}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}A^{T}b\right\|.$
From equation (15), the difference between the LS solution and the HR solution is bounded. As it is known that the LS solution is an unbiased solution for problem (2), the difference between the LS solution and the HR solution can be seen as a bias of the HR solution. The following discussion of the approximation residual of the inverse $(A^{T}A)^{-1}$ and equation (14) show that this bias can be calculated for any given $A$ before the estimation of the solution of the problem (2), which is very important for some problems, such as mobile robot localization problems that estimate the solutions with the same given $A$ but with different $b$. From equation (17), the difference between the TR solution and the HR solution is bounded.
For an HR solution, if $k>0$, then according to the triangle inequality of a norm, the error bound of the HR solution is smaller than the error bound of the TR solution. In other words, the HR solution is closer to the unbiased solution of the ill-conditioned problem under the least squares criterion. And from equation (9), it is achieved by approximating the inverse of the matrix $A^{T}A$, which means that the over-smoothing problem in the TR is solved in the sense of matrix inverse approximation.
III-D Stability
For the matrix $R\left(A^{T}A+R\right)^{-1}$, if its spectral radius satisfies $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$, then one has
$\lim_{i\rightarrow\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}\rightarrow O_{n}$ (18)
where $O_{n}\in\mathbb{R}^{n\times n}$ is a zero matrix.
For the $(k+1)$th-order solution $\hat{x}^{k+1}_{hr}$ and the $k$th-order solution $\hat{x}^{k}_{hr}$, one has
$\hat{x}^{k+1}_{hr}-\hat{x}^{k}_{hr}=\left(A^{T}A+R\right)^{-1}\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}A^{T}b$ (19)
which follows
$\lim_{k\rightarrow\infty}\left(\hat{x}^{k+1}_{hr}-\hat{x}^{k}_{hr}\right)\rightarrow O$ (20)
where $O$ is the zero vector.
Hence, the HR solution is stable. In the next subsection, we will show that, in some cases, the approximation residual of the matrix inverse has a lower bound and an upper bound.
III-E The approximation residual of the matrix inverse
From the approximation expression (9) and equation (13), the approximation residual of the inverse of the matrix $A^{T}A$ is $F(R(A^{T}A+R)^{-1})$ We will show that there is a relationship of the approximation residual
$F\left(R\left(A^{T}A+R\right)^{-1}\right)\preceq o^{\alpha}\left(R\left(A^{T}A+R\right)^{-1}\right),$ (21)
where $\alpha$ is a constant, $o^{\alpha}\left(R\left(A^{T}A+R\right)^{-1}\right)$ is some matrix of $\alpha$ power of the matrix $R\left(A^{T}A+R\right)^{-1}$, the symbol $\preceq$ represents a generalized inequality relationship (partial ordering) succeeds or equal to of two matrices [51], which is similar to the sum of an infinite geometric series.
Lemma 3.
If the spectral radius $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$, then the approximation residual $F\left(R\left(A^{T}A+R\right)^{-1}\right)$ from expression (9) which is defined in equation (13), can be described by two commuting matrices, and the residual can be given by
$F\left(R\left(A^{T}A+R\right)^{-1}\right)$ (22)
$=\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1},$
and if the matrix $R\left(A^{T}A+R\right)^{-1}$ is diagonalizable, then these two commuting matrices are simultaneous diagonalizable matrices.
Proof.
Rewrite the residual of the approximate expression, then one gets
$F\left(R\left(A^{T}A+R\right)^{-1}\right)=\sum_{i=k+1}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}$ (23)
$=\left(I-\left(R\left(A^{T}A+R\right)^{-1}\right)\right)^{-1}\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}$
$=\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}\left(I-\left(R\left(A^{T}A+R\right)^{-1}\right)\right)^{-1}$
which shows that the two factors $\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}$ and $(R(A^{T}A+R)^{-1})^{k+1}$ in equation (22) commute. If $R\left(A^{T}A+R\right)^{-1}$ is diagonalizable, then these two diagonalizable matrices can be diagonalized simultaneously by some nonsingular matrix. ∎
Theorem 1.
If the matrix $R\left(A^{T}A+R\right)^{-1}$ is symmetric, then the approximation residual shown in equation (22) has an upper bound and a lower bound given as follows
$\frac{1}{1-\lambda_{\min}}o^{k+1}\left(R\left(A^{T}A+R\right)^{-1}\right)\preceq F\left(R\left(A^{T}A+R\right)^{-1}\right)$ (24)
$\preceq\frac{1}{1-\lambda_{\max}}o^{k+1}\left(R\left(A^{T}A+R\right)^{-1}\right)$
where $\lambda_{\min}$ and $\lambda_{\max}$ are the minimum and the maximum eigenvalues of the matrix $R\left(A^{T}A+R\right)^{-1}$, and there is always some $R$ such that the approximation method (9) and the equation (24) can be satisfied.
Proof.
If $R\left(A^{T}A+R\right)^{-1}$ is symmetric, then $R\left(A^{T}A+R\right)^{-1}$ and $\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}$ are symmetric and diagonalizable. Apply the simultaneous diagonalization to the two matrices on the right-hand side of equation (22), and let
$R\left(A^{T}A+R\right)^{-1}=P\operatorname{diag}\left\{\lambda_{\max},\ldots,\lambda_{\min}\right\}P^{-1},$ (25)
then one can get
$\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}$ (26)
$=P\operatorname{diag}\left\{\frac{1}{1-\lambda_{\max}},\ldots,\frac{1}{1-\lambda_{\min}}\right\}P^{-1}.$
Note that $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$, then one has
$\frac{1}{1-\lambda_{\min}}I\preceq\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}\preceq\frac{1}{1-\lambda_{\max}}I.$ (27)
Multiplying $\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}$ on all sides of equation (27), then one obtain equation (24). From equation (11), there always exist some regularization matrices that meet the spectral radius condition and ensure that the matrix $R\left(A^{T}A+R\right)^{-1}$ is symmetric. ∎
It can be found that, for $R\in\mathrm{S}_{+}^{n},A^{T}A\in\mathrm{S}_{++}^{n}$, the new SPD matrix $(A^{T}A+R)$ is a convex function with respect to the regularization matrix $R$, and if $\left(R\left(A^{T}A+R\right)^{-1}\right)\in\mathrm{S}_{+}^{n}$, then $o^{k+1}\left(R\left(A^{T}A+R\right)^{-1}\right)$ is a convex function with respect to the matrix $R\left(A^{T}A+R\right)^{-1}$. Note that $(A^{T}A+R)$ and $R\left(A^{T}A+R\right)^{-1}$ are simultaneously diagonalizable via congruence by some nonsingular matrix. By some mild relaxations[6], it is possible to derive equivalent convex problems to optimization problems constructed by these simultaneously diagonalizable constraints [52], which is easy to solve.
As the approximation matrix inverse is used to approximate the inverse of $\left(A^{T}A\right)^{-1}$ in a regularization solution, the upper bound in Theorem 1 indicates that the approximation residual is bounded for the regularization solution, while the lower bound shows a calculable bias of the regularization solution. The bounds here also suggest that, for a regularization method, the difference between the approximate problem (5) and the original problem (2) is finite and can be calculated.
III-F Relaxation of the regularization matrix in the high-order regularization method
To determine the regularization matrix $R$, similar to the truncated singular value decomposition (TSVD) method, which can be found in [6], one way to improve the performance of the HR method can be choosing some regularization matrix $R$ that only modifies the smallest eigenvalue (or some small eigenvalues) of $(A^{T}A)$. One can take the regularization matrix as
$R=P\operatorname{diag}\left\{\lambda_{R,1},\ldots,\lambda_{R,n}\right\}P^{-1}\in\mathbb{R}^{n\times n}$ (28)
where $\lambda_{R,i},i=1,\ldots,n$ is eigenvalues of the regularization matrix $R$, and $P$ is a unitary matrix whose columns are the eigenvectors of matrix $A^{T}A$.
Theorem 2.
For a regularization matrix defined in (28), if $\lambda_{R,i}=\max\left\{\mu^{2}-\lambda_{i},0\right\},i=1,\ldots,n$ where $10<\lambda_{i+1}\leq\lambda_{i},i=1,\ldots,n-$ 1 are the eigenvalues of the matrix $A^{T}A$, and $\lambda_{s+1}\leq\mu^{2}\leq\lambda_{s},0\leq s\leq n-1$, then the regularization matrix only modifies some small eigenvalues of $A^{T}A$ that are smaller than $\lambda_{s}$; and if $\mu^{2}=\lambda_{s}$, the new matrix $A^{T}A+R$ can be given as
$A^{T}A+R=P\operatorname{diag}\left\{\lambda_{1},\ldots,\lambda_{s-1},\lambda_{s},\ldots,\lambda_{s}\right\}P^{-1}$ (29)
where $A^{T}A+R$ is a positive definite matrix, $P$ is a unitary matrix whose columns are the corresponding eigenvectors of matrix $A^{T}A$, and the regularization matrix $R$ of equation (29) satisfies the spectral radius condition.
Proof.
From equations (28) and (29), the given $R$ and $A^{T}A$ commute, so they can be diagonalized simultaneously by some unitary matrix $P$. Let
$A^{T}A=P\operatorname{diag}\left\{\lambda_{1},\ldots,\lambda_{s},\lambda_{s+1},\ldots,\lambda_{n}\right\}P^{-1}$ (30)
and
$R=P\operatorname{diag}\left\{\max\left\{\mu^{2}-\lambda_{1},0\right\},\ldots,\max\left\{\mu^{2}-\lambda_{n},0\right\}\right\}P^{-1},$ (31)
note that $\lambda_{s+1}\leq\mu^{2}\leq\lambda_{s}$, then one can get the new SPD matrix
$A^{T}A+R=P\operatorname{diag}\left\{\lambda_{1},\ldots,\lambda_{s},\mu^{2},\ldots,\mu^{2}\right\}P^{-1}.$ (32)
Note that $0<\lambda_{i}$ and $\lambda_{s+1}\leq\mu^{2}\leq\lambda_{s}$, so the condition $\rho\left(R\left(A^{T}A+R\right)^{-1}\right)<1$ will always be satisfied.
When choosing $\mu^{2}=\lambda_{s}$, then complete the proof. ∎
As a consequence of the Theorem 2, the high-order regularization improves the numerical stability of the solution for ill-conditioned problems when the regularization matrix is chosen according to the equation (29). The numerical stability is enhanced as the condition number of the new matrix $A^{T}A+R$ is smaller than the condition number of the matrix $A^{T}A$ when ${\lambda_{s}}>{\lambda_{n}}$, and the new matrix $A^{T}A+R$ has a condition number $\frac{\lambda_{1}}{\lambda_{s}}$ with respect to the spectral norm while the condition number of the matrix $A^{T}A$ is $\frac{\lambda_{1}}{\lambda_{n}}$. The regularization matrix $R$ is a function of the variables $\mu$ and $s$, hence if $s$ can be determined then the regularization matrix $R$ is a function only with respect to $\mu$. For many robot localization problems, $s$ is not larger than 3 and is not difficult to determine.
This relaxation (28) of $R$ provides a more explainable and reasonable regularization for the ill-conditioned problem as it only changes some small eigenvalues (or the smallest eigenvalue) of the matrix $A^{T}A$. Some methods [6, 7, 10] for improving the SVD method can be used to relax the regularization matrix in the proposed high-order regularization method after minor modifications. The regularization matrix under the relaxation (28) makes sure the conditions in the previous theorems of this work are satisfied. However, replacing the small eigenvalues with some $\lambda_{s}$ directly may lead to losing the information on dimensions with respect to these small eigenvalues. Therefore, it is better to consider a balance between modifying the small eigenvalues and keeping the information of some dimensions with respect to these small eigenvalues while choosing a regularization matrix.
III-G A priori criterion to improve the numerical stability of the ill-conditioned problem
We propose a priori criterion to improve the numerical stability of the solution of problem (2) associated with a matrix $A$, which leads to a numerical stable behavior of the proposed method. To obtain the optimal regularization matrix, we introduce a priori criterion with its objective function given in equation (7).
The first term of the objective function is the residual of approximating the inverse of $A^{T}A$, which is the reason for the estimation bias of the high-order regularization method, and the second term is the condition number of the new matrix $A^{T}A+R$, which describes the numerical stability of the problem. And note that, if $R\in\mathrm{S}_{+}^{n}$, then the new matrix $A^{T}A+R$ is an SPD matrix, and according to Theorem 2, the spectral radius condition can always be satisfied. The proposed criterion considers the trade-off between modifying the small eigenvalues and keeping the information of some dimensions with respect to these small eigenvalues.
Instead of calculating the loss function directly, one can modify the loss function (7) slightly by adding some reasonable relaxations on the regularization matrix $R$ according to the discussion in Subsection (III-F) and solve the loss function more easily. The simplification loss function is
$\min_{s,k\in\mathbb{N},\mu\in\mathbb{R},R\in\mathrm{S}_{+}^{n}}f_{1}(s,k,\mu,R)$ (33)
$=\min_{s,k\in\mathbb{N},\mu\in\mathbb{R},R\in\mathrm{S}_{+}^{n}}\Bigg{\{}\left\|\sum_{i=k+1}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}\right\|$
$\quad\quad\quad\quad\quad\quad\quad\quad+\left\|\left(A^{T}A+R\right)\right\|\left\|\left(A^{T}A+R\right)^{-1}\right\|\Bigg{\}},$
[…]
where the regularization matrix has a special constraint. From equation (29), under these constraints of $R$, one can allow the regularization matrix to only modify some small eigenvalues. In general, if considering only modifying the smallest eigenvalue of the matrix $A^{T}A$, one can calculate the loss function by setting $s=n-1,k=1$. And note that the regularization matrix $R$ is related to $\mu^{2}$, so only $\mu^{2}$ needs to be determined among the parameters in the prior criterion.
Theorem 3.
For $s=n-1,k=1$, the loss function described in equation (33) can be simplified as
$\min_{\mu^{2}\in\mathbb{R}}f\left(\mu^{2}\right)=\min_{\mu^{2}\in\mathbb{R}}\left\{\frac{\left(\mu^{2}-\lambda_{n}\right)^{2}}{\lambda_{n}\mu^{2}}+\frac{\lambda_{1}}{\mu^{2}}\right\},$ (34)
[…]
where $f\left(\mu^{2}\right)$ is a convex function with respect to $\mu^{2}$, and the loss function (34) has a unique minimum at
$\mu^{2}=\min\left\{\sqrt{\lambda_{n}^{2}+\lambda_{n}\lambda_{1}},\lambda_{n-1}\right\}.$ (35)
Proof.
From equation (33), if $s=n-1$, one can get
$\left(A^{T}A+R\right)$ $=P\operatorname{diag}\left\{\tilde{\lambda}_{1},\ldots,\tilde{\lambda}_{n}\right\}P^{-1}$ (36)
$=P\operatorname{diag}\left\{\lambda_{1},\ldots,\lambda_{n-1},\mu^{2}\right\}P^{-1}$
where $\tilde{\lambda}_{i}=\max\left\{\lambda_{i},\mu^{2}\right\},i=1,\ldots,n$ is the eigenvalue of the matrix $\left(A^{T}A+R\right),0<\tilde{\lambda}_{i}$, and $P$ is some unitary matrix.
The condition number of the special new SPD matrix $A^{T}A+R$ is
$\kappa\left(A^{T}A+R\right)$ $=\left\|\left(A^{T}A+R\right)\right\|\left\|\left(A^{T}A+R\right)^{-1}\right\|$ (37)
$=\frac{\tilde{\lambda}_{1}}{\tilde{\lambda}_{n}}=\frac{\lambda_{1}}{\mu^{2}}.$
For the residual, if $k=1$, one can get
$\sum_{i=2}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}$ (38)
$=\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}\left(R\left(A^{T}A+R\right)^{-1}\right)^{2}.$
For the factors in the first term of the equation (33), the matrix power series can be rewritten as
$R\left(A^{T}A+R\right)^{-1}=P\left(\operatorname{diag}\left\{0,\ldots,0,\frac{\mu^{2}-\lambda_{n}}{\mu^{2}}\right\}\right)P^{-1},$ (39)
and one has
$\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}=P\left(\operatorname{diag}\left\{1,\ldots,1,\frac{\mu^{2}}{\lambda_{n}}\right\}\right)P^{-1}.$ (40)
Hence, for the first term in the loss function (33), one can get
$\left\|\sum_{i=2}^{\infty}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}\right\|=\frac{\left(\mu^{2}-\lambda_{n}\right)^{2}}{\lambda_{n}\mu^{2}}.$ (41)
Substitute equations (37) and (41) into the equation (33), then one can get the loss function (34).
Consider the second derivative of $f\left(\mu^{2}\right)$, then it follows that $f\left(\mu^{2}\right)$ is a convex function with respect to $\mu^{2}$. Hence, for the loss function (34), there must be a unique minimum for this function. It can be found that
$\mu^{2}=\min\left\{\sqrt{\lambda_{n}^{2}+\lambda_{n}\lambda_{1}},\lambda_{n-1}\right\}$ (42)
which is the solution for equation (34). ∎
Similarly, for $s=n-1,k=0$, the loss function (33) can be simplified as
$\min_{\mu^{2}\in\mathbb{R}}f\left(\mu^{2}\right)=\min_{\mu^{2}\in\mathbb{R}}\left\{\frac{\mu^{2}-\lambda_{n}}{\lambda_{n}}+\frac{\lambda_{1}}{\mu^{2}}\right\},$ (43)
[…]
Then the corresponding optimal regularization parameter is
$\mu^{2}=\left\{\begin{array}[]{ll}\min\left\{\sqrt{\frac{2\lambda_{1}}{\lambda_{n}}},\lambda_{n-1}\right\},&\text{if }\sqrt{\frac{2\lambda_{1}}{\lambda_{n}}}\geq\lambda_{n}\\ \max\left\{\sqrt{\frac{2\lambda_{1}}{\lambda_{n}}},\lambda_{n}\right\},&\text{if }\sqrt{\frac{2\lambda_{1}}{\lambda_{n}}}\leq\lambda_{n-1}\end{array}\right.$ (44)
which provides the optimal regularization matrix $R$ of the TR method under the proposed a priori criterion.
Additionally, for any given $k\in\mathbb{N},s=n-1$, the loss function is a convex function with respect to $\mu^{2}$, then there must be a unique minimum for the loss function (33), and the optimal regularization matrix can always be found (see Appendix B).
For robot localization problems, Theorem 3 provides a direct way to calculate the regularization matrix $R$ rather than an empirical method in most applications of regularization methods.
III-H Improved regularization solution
In this section, we propose two techniques to improve the proposed HR method for some applications in robot problems. In most robot applications, such as mobile robot localization [43, 44] and calibration of LiDAR-IMU Systems [53], the estimation of the solution includes a multiple estimation process using a number of measurements. It is known that the LS solution is an unbiased estimator for variables with zero mean white noise, and from equation (14), the difference between LS solution and the high-order solution can be calculated in every calculation. Hence, one can calculate the bias of the HR method by calculating the difference between the LS solution and the HR solution. As matrix $A^{T}A$ is nonsingular, and if one has $l$ measurements of $b$, which are a significant amount of the data with zero mean white noise, then one can calculate the bias of the HR method as
$\Delta x$ $=\frac{1}{l}\sum_{t=1}^{l}\left(\hat{x}_{hr,t}-x_{t}^{*}\right)$ (45)
$\approx\frac{1}{l}\sum_{t=1}^{l}\left(\hat{x}_{hr,t}-x_{t}^{*}\right)-\frac{1}{l}\sum_{t=1}^{l}\left(\hat{x}_{ls,t}-x_{t}^{*}\right)$
$=\frac{1}{l}\sum_{t=1}^{l}\left(\hat{x}_{hr,t}-\hat{x}_{ls,t}\right)$
where $\hat{x}_{hr,t}$ is the $t$ th calculation result of the HR method using the $t$ th measurement of $b,\hat{x}_{ls,t}$ is the $t$ th calculation result of the LS method, and $x^{*}$ is the true value. The expression works because the LS solution is an unbiased estimator for a random variable with zero mean white noise and the difference between the LS solution and the high-order solution can be calculated. And then a bias-corrected HR solution can be given by
$\hat{x}^{k}_{hr,l}=\left(A^{T}A+R\right)^{-1}\sum_{i=0}^{k}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}A^{T}b-\Delta x.$ (46)
One equivalent way to calculate the bias of the high-order regularization solution is calculating the approximation residual (22). However, as can be found in equation (40), if the problem is ill-conditioned, then basically, the matrix $\left(I-R\left(A^{T}A+R\right)^{-1}\right)^{-1}$ is ill-conditioned. Hence, it is not able to improve the estimation performance of the high-order regularization method in one calculation. Some similar bias-correction methods can be found in [17].
Let $R\in\mathrm{S}_{+}^{n}$, according to Proposition 1, $R\left(A^{T}A+R\right)^{-1}$ is symmetric positive semidefinite. From equation (24), if the matrix $R\left(A^{T}A+R\right)^{-1}$ is symmetric, then there is an adjustment parameter $\omega$ such that
$\min_{\lambda_{\min}\leq\omega\leq\lambda_{\max}}\left\|G\left(\omega\right)\right\|_{F}^{2},\text{if }R\in\mathrm{S}_{+}^{n}$ (47)
where
$G\left(\omega\right)$ (48)
$=F\left(R\left(A^{T}A+R\right)^{-1}\right)-\frac{1}{1-\omega}\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}$
which shows that the residual can be approximated by a convex function with respect to the PSD matrix $R\left(A^{T}A+R\right)^{-1}$. Here $\|\cdot\|_{F}$ is the Frobenius norm, and $\lambda_{\max}$ and $\lambda_{\min}$ are the maximum and the minimum eigenvalues of matrix $R\left(A^{T}A+R\right)^{-1}$.
As a possible improvement, if the performance of the high-order regularization method for a given $k$ still needs to be improved, from equation (24), one way to improve it can be calculating the following solution
$\hat{x}^{k}_{hr,\omega}=$ (49)
$\left(A^{T}A+R\right)^{-1}\sum_{i=0}^{k}\left(R\left(A^{T}A+R\right)^{-1}\right)^{i}A^{T}b$
$+\frac{1}{1-\omega}\left(A^{T}A+R\right)^{-1}\left(R\left(A^{T}A+R\right)^{-1}\right)^{k+1}A^{T}b,$
$0\leq\omega\leq\lambda_{\max}$
which is an improved solution of the high-order solution. And for $k=0$, one can get
$\hat{x}^{0}_{hr,\omega}=\frac{1}{1-\omega}\left(A^{T}A+R\right)^{-1}A^{T}b,0\leq\omega\leq\lambda_{\max}$ (50)
where $\omega$ is an adjustment parameter based on the performance of the $k$th-order solution. To determine the adjustment parameter, some experimental approaches introduced in [6, 13] can be helpful. There are some researches that introduce a similar adjustment parameter for each diagonal item of the $A^{T}A$ [7, 10]. However, for robot localization problems, if the performance of the high-order regularization is sufficient, for the sake of simplicity, the suggestion adjustment parameter is the minimum eigenvalue $\lambda_{\min}$ or 0.
IV Robot localization in 3D environment
The restricted mobile robot localization problems using sensor networks are often found to be ill-conditioned inverse problems, especially when the trajectory of the robot is random and the placement of the sensors is limited. A mobile robot localization problem can be formulated into the form as (1) with
$A$ $=\left[\begin{array}[]{c}\left(p_{1}-p_{m+1}\right)^{T}\\ \vdots\\ \left(p_{m}-p_{m+1}\right)^{T}\end{array}\right],$ (51)
$b$ $=\frac{1}{2}\left[\begin{array}[]{c}p_{1}^{T}p_{1}-p_{m+1}^{T}p_{m+1}+d_{m+1}^{2}-d_{1}^{2}\\ \vdots\\ p_{m}^{T}p_{m}-p_{m+1}^{T}p_{m+1}+d_{m+1}^{2}-d_{m}^{2}\end{array}\right]$
where $x=\left(x_{r},y_{r},z_{r}\right)^{T}$ is the position of the mobile robot, $A\in\mathbb{R}^{m\times 3},p_{i}=\left(x_{i},y_{i},z_{i}\right)^{T},i=1,\ldots,m+1$ are the positions of $m+1$ anchors with known given positions, and $d_{i}$ are the distance measurements between the $i$ th anchor and the robot. Note that the normal matrix $A^{T}A$ can be expressed as
$A^{T}A$ $=\sum_{i=1}^{m}\left(p_{i}-p_{m+1}\right)\left(p_{i}-p_{m+1}\right)^{T}$ (52)
$=\left[\begin{array}[]{ccc}\Delta x_{r}^{T}\Delta x_{r}\quad\Delta x_{r}^{T}\Delta y_{r}\quad\Delta x_{r}^{T}\Delta z_{r}\\ \Delta y_{r}^{T}\Delta x_{r}\quad\Delta y_{r}^{T}\Delta y_{r}\quad\Delta y_{r}^{T}\Delta z_{r}\\ \Delta z_{r}^{T}\Delta x_{r}\quad\Delta z_{r}^{T}\Delta y_{r}\quad\Delta z_{r}^{T}\Delta z_{r}\end{array}\right]$
where
$\Delta x_{r}=\left(x_{1}-x_{m+1},\quad\dots,\quad x_{m}-x_{m+1}\right)^{T},$ (53)
$\Delta y_{r}=\left(y_{1}-y_{m+1},\quad\dots,\quad y_{m}-y_{m+1}\right)^{T},$
$\Delta z_{r}=\left(z_{1}-z_{m+1},\quad\dots,\quad z_{m}-x_{m+1}\right)^{T}$
and $A=\left[\Delta x_{r},\Delta y_{r},\Delta z_{r}\right]$. From equation (52), if some column of $A$ is much smaller than the other columns, for example, $\Delta z_{r}$ is much smaller than the other two columns, there are some small entries in $A^{T}A$, which causes the smallest eigenvalue of $A^{T}A$. This situation often arises in indoor robot localization problems as the height of a room and the width of a narrow corridor in a building are fixed and much smaller than the value in the other dimensions. In these cases, the localization problems become ill-conditioned problems. Therefore, the position estimation of the robot should be calculated by the high-order regularization method.
To obtain the bias of the HR solutions, one way is to calculate the mean of the difference between every HR solution and LS solution from the start of the estimation process $t=t_{0}$ to the current instant $t=t_{c}$. However, to make sure the estimation of the bias of HR solutions is as accurate as possible, we propose an algorithm based on a sliding window to calculate the difference between the HR solutions and LS solutions. By using the sliding window $W$, which includes $\Delta x(t),t=t_{c}-l+1,...,t_{c}$, we can only consider the mean of differences $\Delta x(t)$ between $l$ HR solutions and $l$ LS solutions from instant $t=t_{c}-l+1$ to the current instant $t=t_{c}$. Therefore, we only need to consider the latest $l$ measurements of $b$ and assume that in this time interval, the error of $b$ is a zero mean white noise and $l$ large enough. The biased corrected algorithm based on a sliding window is given as follows.
Algorithm 1 Bias-correction of the HR solution with a sliding window.
  Bias sliding window$(W)$
       if $\textbf{if }t_{c}-t_{0}<l$
       $W\leftarrow\Delta x(t)=\hat{x}_{hr,t}-\hat{x}_{ls,t}$ for $t=t_{0},...,t_{c}$
       $\Delta x=0$
       else if $\textbf{else if }t_{c}-t_{0}\geq l$
       $W\leftarrow\Delta x(t)=\hat{x}_{hr,t}-\hat{x}_{ls,t}$ for $t=t_{c}-l+1,...,t_{c}$
       $\Delta x=\frac{1}{l}\sum\left(\Delta x(t)\right),\Delta x(t)\in W$
       return $\Delta x$
  HR Solution$(\hat{x}_{hr})$
       $\hat{x}_{hr}\leftarrow\hat{x}_{hr}-\Delta x$
       return $\hat{x}_{hr}$
The robot localization process is shown in Fig. 1. In Fig. 1, a robot moves continuously along a route. The distances between anchors and the robot are given by a tag, which is a distance measurement sensor equipped with the robot. The positions of the robot in different instants can be obtained by using distance measurements.
Figure 1: Illustration of robot localization problem using anchors.
V Simulations and Experiments
To demonstrate the performance of the proposed methods, we compare the results of the high-order method (HRM) with the LS method, the TR method, Fuhry’s Tikhonov regularization (FTR) method with the optimal regular parameter (OFTR), which is a special case of HR with $\mu$ provided by the equation (44). As the FTR bridges the TR and the TSVD with a filter factor[6], the relationships between TSVD and the proposed HR can also be found through such a similar filter factor. In addition, many of the modified Tikhonov regularizations[7, 10, 8, 9] use a similar form to the FTR only with a different filter factor. Hence, the relationships between these modified Tikhonov regularizations and the proposed HRM can be discovered. Consider $s=n-1$, some special cases of the proposed HRM are shown in Table I. As these special cases and TSVD are all the fundamental regularization methods and more importantly their regularization matrix or regularization parameters can be determined by a priori criterion, i.e., existing a closed-form solution, we then compare the performance of the proposed HRM with such regularization methods. We also provide the optimal regularization parameter $\mu$ of the proposed HRM in the special case for $k=1$ in Table I, which is used to compare with the other regularization methods in simulations and experiments.
TABLE I: Some special cases of the proposed HRM
Method $k$ $R$ $\omega$
LS Inf 0 -
TR 0 $R=\mu^{2}I$ 0
FTR 0 $\lambda_{R,n}=\max\left\{\mu^{2}-\lambda_{n},0\right\}$ 0
OFTR 0 $\mu^{2}=(2\lambda_{1}/\lambda_{n})^{1/2}$ 0
HR $>0$ $\mu^{2}=(\lambda_{n}^{2}+\lambda_{n}\lambda_{1})^{1/2}(k=1)$ 0 or $\lambda_{min}$
V-A Simulations
We first create a 3D scenario for random point localization to show the performance of the proposed high-order method. In these simulations, we create 1000 points randomly representing the robot positions. We add zero mean random noise with $\sigma^{2}=0.1m$ to each distance measurement according to the results in [54], and then each entry of $b$ is with random noise. For this 3D localization problem, the results shown in Table II, Table III and Fig. 2 indicate that the performance of the proposed HR method is significantly spurior than the LS method, the TR method, and Fuhry’s Tikhonov regularization (FTR) method with the optimal regular parameter (OFTR), which is provided by the equation (44).
TABLE II: Results of different methods (unit: meter)
Method LS TR HRω HRλ
RMSE 0.64331 0.42688 0.39173 0.32005
Figure 2: The performance of the high-order method with $R$ only changing the smallest eigenvalue of $A^{T}A$ and $\mu^{2}=(\lambda_{n}^{2}+\lambda_{n}\lambda_{1})^{1/2}(s=n-1,k=1)$, the OFTR method and the LS method. The bias of the HR solution is corrected in real-time according to the LS solution. The result of the OFTR solution is provided by the FTR method whose regularization parameter $\mu^{2}=(2\lambda_{1}/\lambda_{n})^{1/2}$ is given by the a priori criterion proposed in this paper for $s=n-1,k=0$, which only changes the smallest eigenvalue of the matrix $A^{T}A$. (a) Total RMSE. (b) RMSE in different dimensions.
TABLE III: Results of the OFTR method and the HR method with optimal regularization parameter
Method Total RMSE RMSE in z
LS 0.64654 0.62970
OFTR($\mu^{2}=(2\lambda_{1}/\lambda_{n})^{1/2}$) 0.29959 0.25928
HR($\mu^{2}=(\lambda_{n}^{2}+\lambda_{n}\lambda_{1})^{1/2}$) 0.27175 0.22592
In Table II, the TR method is with the regularization matrix $R=\mu^{2}I$, the HR method with the same regularization matrix and an adjustment parameter $\omega=\lambda_{min}$ is denoted as HRω, and the HR method with only changing the smallest eigenvalue of $A^{T}A$ with $\lambda_{R,n}=\lambda_{n-1}$ is denoted as HRλ. From Table II, the adjustment parameter and the relaxation of the regularization matrix help to improve the behavior of regularization. It can be known from the simulation results of Fig. 2 and Table III that the high-order regularization method is superior to the OFTR method, which is the special case of HR method for $k=0$, indicating the order $k$ of regularization benefits the overall behaviors of regularization. The estimation error in $z$ dimension also shows that the high-order regularization method with the proposed a priori criterion finds a balance between modifying the small eigenvalues and keeping the information of the dimension with respect to the smallest eigenvalue as its performance in the $z$ dimension is improved.
To show the performance improvement of the high-order regularization method in the dimension with respect to the smallest eigenvalue of $A^{T}A$ and the trade-off capability of the proposed optimal regularization matrix for the estimation bias, over-smoothness and the numerical stability, we create a 3D scenario in which the robot moves along a given route shown in Fig. 3. In this scenario, the robot moves along the given route which consists of two main parts: in the first part, the robot moves on a horizontal plane, and in the second part of the movement, the robot moves on a slope. We also add zero mean random noise with $\sigma^{2}=0.1m$ to each distance measurement as in the previous simulations.
Figure 3: A given route to simulate the robot moves in a real 3D environment.
In order to show the over-smoothness of some regularization solutions (such as TSVD) and show the performance of the HR method after the bias correction, we perform another simulation based on Algorithm 1. The results of different methods are shown in Fig. 4 and Table IV. In this experiment, the bias of the HR solution is corrected by the mean of the difference between the HR solution and the LS solution based on a sliding window.
Figure 4: The performance of the high-order method with $R$ only changing the smallest eigenvalue of $A^{T}A$ and $\mu^{2}=\lambda_{n-1}(s=n-1,k=1)$, the TSVD method and LS method. The bias of the HR solution is corrected in real-time based on Algorithm 1 by using the latest $l$ measurements with a sliding window. And TSVD only truncates the smallest singular value with respect to the z component. In this simulation, the robot moves along a given route. (a) Total RMSE. (b) RMSE in different dimensions.
TABLE IV: Results of different methods and the bias-corrected HR solution based on Algorithm 1
Method Total RMSE RMSE in x RMSE in y RMSE in z
LS 0.61150 0.09077 0.11994 0.59271
TSVD 0.31743 0.09905 0.12209 0.27576
HR 0.19921 0.09525 0.11857 0.12865
From Fig. 4 and the numerical result (Table IV), the RMSE of the HR solution in $z$ dimension is almost the same with the RMSE in $y$ and $x$ dimension, which means that, for such an ill-conditioned robot localization problem, the HR method gives a numerically stable solution just as it does for the well-conditioned problems. In comparison, even though the result of the TSVD performs better than the LS method, the over-smoothness of the TSVD is severe, which can be easily found on the error in $z$ dimension from Fig. 4. This result again indicates the excellent trade-off capability of the proposed high-order regularization method for the estimation bias, over-smoothness and numerical stability. The results in Table IV reveal that the proposed HR method markedly improves localization accuracy, as evidenced by the total RMSE and the RMSE in the $z$ dimension, which shows enhancements of $67.4\%$ and $78.3\%$, respectively. These results are obtained under stringent conditions where the anchor heights are highly limited, with a maximum height differential of 0.5 meters in the simulations.
V-B Experiments
In this section, we describe an extensive experimental validation in the real-world environment to evaluate the performance of the proposed method. Our experimental platform is shown in Fig. 5. Four anchors are used in these experiments to locate the robot by a tag carried on this robot. The coordinates of the anchors are measured manually using a rangefinder and the height of one anchor is changed in different experiments to reach the different degrees of ill-conditioned situations. The positions of the four anchors are (0, 0, 2.29), (5.30, 4.12, 1.20), (-0.56, 2.01, 0.30), and (6.00, 0, 1.20) ((6.00, 0, 1.80) in another experiment). The robot is designed to move along a given trajectory with unknown dynamics. The position of the robot is calculated only by the current distance measurements between the robot and the four anchors, which means that no assumption of the robot dynamics and the trajectory is required to calculate the position of the robot. An Ouster Lidar is used to provide the reference positions of the robot, which are used to evaluate the performance of the proposed method in this paper.
Figure 5: The platform in our experiment. The white box on top is the UWB sensor that measures the distances between the robot and the anchors, and the Lidar under the white box is used to provide the reference positions of the robot
In order to reduce the influence of some unexpected factors such as environment disturbance and UWB sensor antenna orientation on distance measurements, a filter is applied to filter the data before calculating the position of the robot. To estimate the performance of the proposed method, the Umeyama method[55] is used in these experiments to align two point patterns into the same frame referring to some evaluation tools[56]. The result and the reference of the localization experiment are shown in Fig. 6.
Figure 6: The estimation result (est) and the reference path (ref) of the robot moving in a real 3D indoor environment. The reference path is calculated using the LiDAR measurements to compare the estimation of the proposed HR method.
Figure 7: The experiment results. The result is given by the high-order method with $R$ only changing the smallest eigenvalue of $A^{T}A$ and $\mu^{2}=(\lambda_{n}^{2}+\lambda_{n}\lambda_{1})^{1/2}(s=n-1,k=1)$. The result of the OFTR solution is provided by the FTR method whose regularization parameter $\mu^{2}=(2\lambda_{1}/\lambda_{n})^{1/2}$ is given by the a priori criterion proposed in this work for $s=n-1,k=0$, which only changes the smallest eigenvalue of the matrix $A^{T}A$.
Figure 8: The detailed performance of the proposed high-order method in a real 3D environment.
TABLE V: Experimental results of different methods
Method RMSE Mean Max Min
LS 0.28289 0.26380 0.61741 0.07171
TSVD 0.24792 0.22738 0.53397 0.05278
OFTR 0.25128 0.23150 0.54278 0.05656
HR 0.23519 0.21519 0.54860 0.05656
TABLE VI: Experimental results of different methods with a different height of one anchor
Method RMSE Mean Max Min
LS 0.26989 0.25256 0.56059 0.06554
TSVD 0.24516 0.23229 0.48323 0.08341
OFTR 0.24557 0.23317 0.48225 0.08240
HR 0.21222 0.19865 0.45421 0.06630
The experiment results in Fig. 7 indicate that the proposed HR method improves the localization performance in z-dimension which is the same as the simulations. Comparing the results of OFTR (one special case of HR with $k=0$ and optimal $R$) and HR ($k=1$) methods, we find that the error of OFTR in z-dimension is smaller than HR, which is a severely over-smoothing situation, as the error in z-dimension is even much smaller than errors in x-dimension and y-dimension. This situation occurs during our robot moves nearly in a plane. In some similar cases, we can even obtain a smaller error in z-dimension by assuming the change in z-dimension for the robot is 0, and we can get the same small error result by using a larger regularization parameter $\mu$ for regularization matrix $R$. However, the assumption does not hold for many applications (such as for drones) and is not required in the proposed high-order regularization. From Fig. 7, the experiment results of the HR method in different dimensions are almost the same, which means that our high-order method overcame the over-smoothing problem and ensures a high localization performance. We also perform another experiment by changing the height of one anchor to show the behavior of the proposed HR method under different degrees of ill-conditioned situations. The detailed performance of the proposed high-order method is shown in Fig. 8, and all numerical results are given in Table V and Table VI. From the experiments results in Table VI, the proposed HR method demonstrates excellent performance of improvement in RMSE and mean error by over $20\%$ with the best enhancement of $21.3\%$, even when the anchor heights are not significantly limited, as in the experiments, the maximum height variation of anchors is 1.99 meters.
VI Conclusion and discussion
In this work, we propose a high-order regularization method to solve the restricted robot localization problem when the associated matrix is ill-conditioned. From the simulation, in a severely restricted scenario, the proposed HR method archives a significant enhancement of $78.3\%$ in $z$ dimension (the restricted dimension), while in a scenario without severe restrictions, the improvement of the proposed method also reaches more than $20\%$ from the experiment results. We show that the proposed method is superior to many Tikhonov regularization-based methods in approximating the inverse problem. It is shown that the Tikhonov regularization can be seen as a special case of the proposed method. The explanation of the over-smoothing of the Tikhonov regularization solution is also given. The proposed method is based on eigenvalue decomposition, which means that there will always be some explainable and reasonable techniques to improve the estimation performance of the proposed method as eigenvalue decomposition can always be applied to the new matrix.
To find an explainable and reasonable regularization matrix, we relaxed the regularization matrix with a special form, which only changes some small eigenvalues of the matrix $A^{T}A$. We also proposed a priori criterion to improve the numerical stability of the problem, which can be seen as replacing the ill-conditioned problem with a well-conditioned approximate problem. The optimal regularization matrix can be calculated based on the proposed criterion, which considers the balance between modifying small eigenvalues and keeping the information of some dimensions with respect to these small eigenvalues. However, if the information on some dimensions is not needed or can be obtained by some other technologies, the proposed criterion may not provide the best choices for the regularization matrix. The optimal regularization matrix also provides the potential to verify if the given $k$ or $s$ is appropriate. Two methods to improve the performance of the $k$th-order regularization and calculate the bias of the regularization solution are also proposed, which may help achieve a possible unbiased solution similar to the LS solution.
For future work, some possible improvements of the high-order regularization method can be determining the adjustment parameter by calculating the residual directly or calculating the bias with some expressions with respect to the new matrix and the regularization matrix. Extending such a high-order regularization to nonlinear model cases will be the next step of this work. And applying the high-order regularization to some other ill-conditioned problems is also a promising topic, especially, applying the proposed explainable and bounded high-order regularization to machine learning, which uses regularization methods widely to achieve generalization abilities, may drive a more explainable machine learning theory.
Acknowledgments
The authors would like to thank Bayu Jayawardhana for the UWB sensors and Simon Busman for his help in the experiments.
Appendix A Proof of a similar approximate expression
Consider the Cholesky factorization of $A^{T}A$ is $S^{T}S$, then one has
$\left(A^{T}A+R\right)=\left(S^{T}S+R\right).$ (54)
If the spectral radius condition $\rho\left(\left(S^{T}\right)^{-1}R(S)^{-1}\right)<1$ is satisfied, then a special and similar approximate expression of the inverse of matrix $A^{T}A$ can be found as follows
$\left(A^{T}A+R\right)^{-1}$ (55)
$=(S)^{-1}\left[I+\left(S^{T}\right)^{-1}R(S)^{-1}\right]^{-1}\left(S^{T}\right)^{-1}$
$=\left(A^{T}A\right)^{-1}\sum_{i=0}^{\infty}\left(-R\left(A^{T}A\right)^{-1}\right)^{i}$
and the spectral radius becomes $\rho\left(-R\left(A^{T}A\right)^{-1}\right)<1$. Then one has
$\left(A^{T}A\right)^{-1}=$ $\left(A^{T}A+R\right)^{-1}+\left(A^{T}A\right)^{-1}R\left(A^{T}A\right)^{-1}$ (56)
$-\left(A^{T}A\right)^{-1}\sum_{i=2}^{\infty}\left(-R\left(A^{T}A\right)^{-1}\right)^{i}$
and
$-R\left(A^{T}A\right)^{-1}+R\left(A^{T}A+R\right)^{-1}$ (57)
$=-R\left(A^{T}A\right)^{-1}\left(R\left(A^{T}A\right)^{-1}-\sum_{i=2}^{\infty}\left(-R\left(A^{T}A\right)^{-1}\right)^{i}\right).$
When $k=1$, the bias between LS solution (3) and the high-order solution (12) is
$\hat{x}_{ls}-\hat{x}_{hr}=$ $\left(A^{T}A\right)^{-1}A^{T}b-\left(A^{T}A+R\right)^{-1}A^{T}b$ (58)
$-\left(A^{T}A+R\right)^{-1}R\left(A^{T}A+R\right)^{-1}A^{T}b$
Note that, one can get
$\left(A^{T}A\right)^{-1}-\left(A^{T}A+R\right)^{-1}-\left(A^{T}A+R\right)^{-1}R\left(A^{T}A+R\right)^{-1}$ (59)
$=\left(\left(A^{T}A\right)^{-1}-\left(A^{T}A+R\right)^{-1}\right)R\left(A^{T}A\right)^{-1}$
$\quad+\left(A^{T}A+R\right)^{-1}\left(R\left(A^{T}A\right)^{-1}-R\left(A^{T}A+R\right)^{-1}\right)$
$\quad-\left(A^{T}A\right)^{-1}\sum_{i=2}^{\infty}\left(-R\left(A^{T}A\right)^{-1}\right)^{i}$
and then according to (56) and (57), one has the approximation error of the matrix inverse
$err_{ap}((A^{T}A)^{-1})=$ (60)
$\left(A^{T}A\right)^{-1}-\left(A^{T}A+R\right)^{-1}-\left(A^{T}A+R\right)^{-1}R\left(A^{T}A+R\right)^{-1}$
$=\left(A^{T}A+R\right)^{-1}\left(\left(I+R\left(A^{T}A\right)^{-1}\right)^{-1}-I+R\left(A^{T}A\right)^{-1}\right).$
Hence, the bias can be described as
$\hat{x}_{ls}-\hat{x}_{hr}$ (61)
$=\left(A^{T}A+R\right)^{-1}$
×$\quad\times\left(\left(I+R\left(A^{T}A\right)^{-1}\right)^{-1}-I+R\left(A^{T}A\right)^{-1}\right)A^{T}b.$
This result shows that when $k=1$, the bias between LS solution and the high-order solution can be described in different forms under some different approximate expressions of the inverse of matrix $A^{T}A$. However, some approximate expressions may not guarantee a stable solution to the ill-conditioned problem as in these approximate expressions the inverse $(A^{T}A)^{-1}$ still needs to be calculated or only suitable for some $k$.
Appendix B Proof of convexity of the loss function
The following proof shows that the loss function is a convex function, for $s=n-1,k\in\mathbb{N}$. For $s=n-1,k=0$ and $k=1$, we have shown that the loss function in (33) is a convex function. Note that when $\mu^{2}=\lambda_{n}$, then $R=O_{n}$ Asumme $s=n-1,k\geq 2$ and $\mu^{2}>\lambda_{n}$, one can get
$g(\mu):=\frac{\mu}{\lambda_{n}}\left(\frac{\mu-\lambda_{n}}{\mu}\right)^{k+1}=\frac{1}{\lambda_{n}}\mu\left(1-\frac{\lambda_{n}}{\mu}\right)^{k+1},$ (62)
$g^{\prime}(\mu)=\left(\frac{\mu+k\lambda_{n}}{\left(\mu-\lambda_{n}\right)\mu}\right)g(\mu),$ (63)
$g^{\prime\prime}(\mu)=\left(\frac{\left(k+k^{2}\right)\lambda_{n}^{2}}{\left(\left(\mu-\lambda_{n}\right)\mu\right)^{2}}\right)g(\mu)$ (64)
$=\frac{\lambda_{n}\left(k+k^{2}\right)}{\mu^{2}\left(\mu-\lambda_{n}\right)}\left(\frac{\mu-\lambda_{n}}{\mu}\right)^{k}>0.$
And then the second derivative of the loss function in (33) can be expressed as
$f^{\prime\prime}(\mu^{2})=g^{\prime\prime}(\mu^{2})+\frac{2\lambda_{1}}{\mu^{6}}$ (65)
$=\frac{\lambda_{n}\left(k+k^{2}\right)}{\mu^{4}\left(\mu^{2}-\lambda_{n}\right)}\left(\frac{\mu^{2}-\lambda_{n}}{\mu^{2}}\right)^{k}+\frac{2\lambda_{1}}{\mu^{6}}>0.$
Hence, for $s=n-1,k\in\mathbb{N}$, the loss function is a convex function with respect to $\mu^{2}$. Therefore, there must be a unique minimum of the loss function, and an optimal regularization matrix $R$ can be found at the minimum of the loss function.